OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value?

OpenStudy (debbieg):

Just plug in x=5 for the x, and evaluate. y=(5 + 3)2 = ? (I'm assuming that the 2 is a factor, as you have it written, and not an exponent).

OpenStudy (anonymous):

If y = (x + 3)2, and the value of x=5, what is the value? A. 12 B. 16 C. 18 D. 14

OpenStudy (debbieg):

OK, given what I explained above, what do you think the answer is? The choices are not needed here.

OpenStudy (anonymous):

it would be A right

OpenStudy (debbieg):

How did you get A? What is (5 + 3)2 ? What do you do first?

OpenStudy (anonymous):

I got it now thanks